Determine the magnetic feild at the center of the semicircular piece of wire with radius 0.20 m. The current carried by the semi circular piece of wire is 150 A.


1
Expert's answer
2022-03-15T10:34:27-0400

The magnetic field at the center of semicircular loop


"B=\frac{\mu_0 I}{4 r}""B=\frac{4\pi*10^{-7}*150}{4*0.20}=2.4*10^{-4}\:\rm T"